The dichromate ion oxidizes iron (II) to iron (III) in acid solution. Balance the equation:
Cr2O
2- + H3O
+ + Fe
2+ → Cr3+ + Fe
3+ + H2O, by the half-reaction method
1
Expert's answer
2019-07-15T07:04:34-0400

Cr2O72- + H3O+ + Fe2+ = Cr3+ + Fe3+ + H2O


2Cr+6 + 6e = 2Cr+3    6      1

Fe+2 – 1e = Fe+3        1      6


Therefore, the ionic equation has the form:


Cr2O72- + 14H3O+ + 6Fe2+ = 2Cr3+ + 6Fe3+ + 21H2O